OpenStudy (anonymous):

Solve the following equation by completing the square. Follow the steps given at the end of this section. x^2-6x = 0

OpenStudy (mertsj):

1/2 of 6 = 3 and 3^2 = 9 so add 9 to both sides. Now you have: x^2-6x+9 = 9 Factor the left side: (x-3)^2 = 9 Take the square root of both sides and get x-3 = 3 or x - 3 = -3 Solve each equation and get x = 6 or x = 0